# Can AR-Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision Making? A Comparative Study of AR See-Through vs. 2D Minimap

### v1

@@ 4.1.1 Accuracy
4.1.2 AI Reliance
Results from the accuracy analysis indicated that participants relied on AI inappropriately in the AR+AI condition. Quantitative analysis confirms this pattern: we observed a significant difference between Minimap+AI and AR+AI, with participants showing less appropriate reliance on AI in AR+AI (Figure 6a). This finding again contradicts our second hypothesis H2.
Inappropriate reliance on AI can be either over-reliance or under-reliance. To examine which pattern dominated participants’ interactions with AI in AR+AI, we separately analyzed trials of over-reliance and under-reliance. Figure 6b and c depict the average occurrences of over- and under-reliance, respectively, per eight trials.
To interpret these numbers, we compared observed behaviors to a random-selection baseline. With four possible options per trial, random selection would result in an expected average of 0.75 over-reliance instances and 4.25 under-reliance instances per eight trials (see supplementary material). Compared to this baseline, participants in the AR+AI condition displayed fewer instances of under-reliance ($M_{\text{{AR}}}=1.63<M_{\text{random}}=4.25$), but notably more instances of over-reliance ($M_{\text{{AR}}}=1.41>M_{\text{random}}=0.75$). In contrast, participants in the Minimap+AI condition exhibited fewer instances of both under-reliance ($M_{\text{{Minimap}}}=0.84<M_{\text{random}}=4.25$) and over-reliance ($M_{\text{{Minimap}}}=0.72<M_{\text{random}}=0.75$).
These findings indicate that inappropriate reliance in the AR+AI condition was primarily driven by over-reliance, consistent with our earlier accuracy analysis (Sec. 4.1.1).
Figure 6: Results of AI reliance across visualizations. Metrics include (A) appropriate reliance, (B) under-reliance, and (C) over-reliance, comparing Minimap+AI and AR+AI conditions. Random-selection baselines (in red) are indicated.
4.1.3 Response Time
We computed the average response time across conditions and observed significant effects (Figure 5b). Participants responded faster in Minimap+AI than AR+AI ($\Delta_{M}=\text{-0.94},p=\text{0.01}$) and in Minimap+NoAI than AR+NoAI ($\Delta_{M}=\text{-2.00},p<\text{0.001}$). These results contradict our null hypothesis H3.
Response Time Excluding Initial Search. We observed that participants in the AR conditions spent additional time to visually searching the spatially distributed targets—especially those outside their immediate field of view. In contrast, in the Minimap conditions, all targets were presented in a 2D overview, making search overhead negligible. Prior decision-making models (e.g., Simon’s model [71], the drift–diffusion model [61], the leaky competing accumulator model [78]) and empirical studies (e.g., [62, 64]) suggest that actual decision making begins only after sufficient information has been gathered. This motivates us to exclude the initial search overhead in the AR conditions to better compare its decision making phase with the Minimap conditions.
Figure 7: (a) and (b) show the direct and walking distance to the orange target; (a) and (b) also indicate the start and end moment of searching for the blue target.
Following prior visualization-decision work [55], we consider that meaningful decision making start only after at least two options have been identified. Thus, for each trial in the AR conditions, we excluded the initial search time required to discover the first two distinct targets. Specifically, we defined two search segments to exclude: S1-from the start of the trial to the first frame when the first target entered the user’s field of view; and S2-from the frame when the user turned away from the first target (e.g., Figure 7a) to the frame when the second target appeared (e.g., Figure 7b). If a target was already visible at trial start, or if the first and second targets appeared simultaneously, the corresponding segment (S1 or S2) was set to zero. This definition provides a strict lower bound on search time, as participants may still require additional head movement to center the target in view. Two co-authors independently annotated these segments frame by frame from the first-person recordings to ensure inter-rater reliability. Discrepancies were discussed and resolved through consensus.
After subtracting the annotated initial search time from the response time, we observed a different pattern (Figure 5c). Participants in AR+AI spent less time than those in Minimap+AI with a significant effect ($\Delta_{M}=\text{-0.85},p=\text{0.01}$), indicating that, with AI assistance, decisions were made more quickly with the embedded visualization. No significant difference was observed between Minimap+NoAI and AR+NoAI, suggesting that, without AI assistance, participants spent comparable time making decisions across visualizations In summary, although raw response times were longer in AR conditions, a detailed breakdown reveals that this extra time wasn’t spent on the decision making. This leads to interesting implications that we will further discuss in Sec. 5.4.
4.1.4 Post-Trial Pointing Error Rate
[…]
4.1.5 Self-Reported Confidence
4.2 Qualitative Results
We conducted semi-structured post-study interviews to capture participants’ strategies, rationales, and motivations behind their decision-making across all conditions. Participants are referred to as P1–P32. We analyzed the interviews using reflexive thematic analysis [7], identifying themes both inductively (bottom-up) and deductively (top-down). Below, we summarize the key themes that emerged from the analysis.
4.2.1 Diverse Decision Strategies Across Visualizations
Strategies in Both Visualizations: People First vs. Distance First. Participants primarily adopted one of two strategies: (1) prioritizing fewer people in line, then considering walking distance; or (2) prioritizing shorter distance, then checking queue length. In the Minimap conditions, participants were almost evenly split between these two strategies, with roughly half employing a “people first” (15 out of 32) and the other half a “distance first” (14 out of 32). By contrast, in the AR conditions, a slight preference emerged for distance-first strategies (15 out of 32), while fewer participants initially considered queue lengths (11 out of 32). The remaining participants described balancing both factors equally without a dominant criterion.
AR: Unique Heuristics and Motivations. Whereas strategy selection in the Minimap conditions appeared to depend largely on participants’ spatial skills, the AR conditions introduced distinct perceptual heuristics influencing participants to prioritize distance:
  • Egocentric Spatial Imagery. AR see-through’s egocentric view led participants to vividly imagine physical navigation, especially complex pathways with hallway turns or floor changes. For example, P9 noted, “It’s easier to draw or imagine a line from where I am to the coffee machine…” This detailed mental imagination caused them to weigh distance more heavily in their decisions.
  • Visual Proximity Bias. Participants described a misleading visual proximity effect, where certain targets appeared visually closer (Figure 7a) due to the see-through nature of the AR visualizations, even if actual walking paths were longer (Figure 7b). As P8 noted, “I think this way is shorter but it’s actually longer.” This illusion occasionally influenced participants to overweight distance.
  • Occlusion Challenges. Participants encountered difficulties estimating queue lengths due to occlusion in the AR visualizations, prompting reliance on distance estimates instead of people counts.
@@ 4.2.1 Diverse Decision Strategies Across Visualizations
Summary. Compared to the Minimap conditions, participants in the AR conditions showed a stronger preference for spatial reasoning (distance) in their decision-making. This suggests that the embedded AR-see through visualization inherently emphasizes spatial perception, which, however, can be a double-edged sword that introduces perceptual biases and occlusion challenges.
4.2.2 Increased Trust in AI Suggestions with AR See-Through
The visualizations impacted participants’ trusted and relied on AI.
@@ 4.2.2 Increased Trust in AI Suggestions with AR See-Through
Minimap: Make Own Choice First, Then Consult AI. In the Minimap conditions, most participants (22 out of 32) preferred first to form their own judgment and then cross-check against the AI suggestion (e.g., “I use it just to double-check my decision”, P22). Participants reported feeling more confident in their own judgment when using the Minimap visualization because it presented the information required to make informed choices. They typically consulted the AI in cases of uncertainty or indecision: “If I felt stuck between two locations, I would check the AI to break the tie” (P20).
@@
AR: AI as a Starting Anchor. Participants expressed greater trust and reliance on AI suggestions in the AR conditions, frequently using the AI suggestion as an initial reference or anchor for their decisions. Several participants (e.g., P6, P11) mentioned that AI suggestions helped narrow the decision space and enabled quicker comparisons. Participants provided several reasons for trusting AI more with AR:
@@
  • Perceptual and navigation challenges. Participants found distance estimation (e.g., P2), queue length assessment (e.g., P4, P5), and locating navigation paths more effortful in AR (e.g., P12). Thus, AI suggestions offered a useful cognitive shortcut, aligning with findings from prior research on cognitive load and decision-making [3].
@@
  • Unfamiliarity with AR See-Through visualizations. Some participants mentioned that unfamiliar with the AR See-Through visualizations increased their reliance on AI guidance (e.g. P31).
@@
  • Enhanced Realism and Embodiment of AI suggestions. Interestingly, several participants indicated that the realistic, embodied way AI suggestions appeared in the AR See-Through increased their sense of trust and acceptance (e.g., P20, P21).
@@
Summary. compared to the Minimap, the AR visualization led to greater participant trust in and reliance on AI suggestions. Participants attributed this shift to greater perceptual complexity, unfamiliarity with the AR visualizations, and the more realistic presentation of AI.
5 Discussion
@@
5.1 Perceptual Challenges of AR-Embedded Visualizations Can Lead to AI Over-Reliance
Our findings suggest that perceptual challenges of embedded visualizations in AR – such as occlusion, large canvas space, and spatial estimation difficulty – can increase users’ reliance on AI suggestions, even when those suggestions are suboptimal. This over-reliance does not appear to stem from blind trust in AI, but rather from the inherent complexity of perceiving and evaluating spatial options in large-scale, dynamic physical environments. While prior research has discussed AR’s perceptual limitations [18], our study directly connects these challenges to AI reliance patterns in time-sensitive spatial decision-making. This highlights the need for better embedded visualization designs that support spatial reasoning, especially in large, dynamic environments.
Design Implications. Providing AI support in AR environments is not inherently beneficial. It can be neutral or even detrimental when perceptual challenges make it harder to verify AI suggestions. These findings underscore the need to rethink how AI assistance is integrated into AR, rather than directly porting AI designs from desktop or 2D interfaces. Designers should explicitly consider how spatial perception difficulties affect users’ ability to evaluate AI output.
Future Studies. In our study, situational data (e.g., queue length) was visualized using human-shaped glyphs. A natural next step is to compare visual encodings (e.g., glyphs vs. numeric labels) to evaluate how data representation influences AI reliance. Additionally, examining how different levels of spatial complexity (e.g., number of targets, layout density) modulates trust and reliance in AR-embedded visualizations would offer deeper insight into AR-AI interaction design.
5.2 Cognitive Biases Triggered by AR-Embedded Visualizations Can Promote Inappropriate AI Reliance
Beyond perceptual difficulty, our qualitative findings point to cognitive biases that are unique to AR-embedded visualizations. Participants reported visual proximity illusions and increased trust in AI suggestions due to their spatial embodiment in the environment. These effects align with prior research on spatial cognition biases [60] and trust in embodied agents [37].
Although cognitive biases have been studied in traditional visualization contexts [15, 82], their implications in embodied, spatial, and AI-augmented AR systems remain underexplored. As Padilla et al. [54] suggested, when systems are designed with such biases in mind, users can benefit from Type 1 (intuitive) thinking—particularly under time pressure. Our findings highlight a fertile area for interdisciplinary research at the intersection of spatial cognition, embedded visualizations, and human-AI decision-making.
@@ 5.2 Cognitive Biases Triggered by AR-Embedded Visualizations Can Promote Inappropriate AI Reliance
5.3 Aligning the Strengths of AR-Embedded Visualizations with Spatial Tasks
@@
In our study, while the AR see-through did not outperform the 2D minimap on decision accuracy, it significantly reduced post-trial pointing errors, suggesting improved spatial mapping and memory. This distinction suggests that AR-embedded visualizations may be more beneficial in continuous, action-based spatial workflows rather than in isolated decision-making tasks.
@@
In high-stakes domains like emergency response, selecting a correct location is only the first step – what follows is action. Misinterpreting a target’s location after selection can be consequential. For instance, a user might correctly choose an exit from a wall-mounted map but walk the wrong way due to reference-frame confusion. AR-embedded visualizations mitigates this risk by spatially anchoring information in the environment. Thus, rather than emphasizing the weakness of AR-embedded visualizations, our findings illustrate its potential strengths in embodied spatial tasks and highlight when and where AR-embedded visualizations are best applied.
@@
Design Implications. AR-embedded visualizations may be most effective when tightly coupled with tasks involving physical action, rather than static decision-making. Designers should consider the viewer’s movement and real-time orientation when designing the embedded visualizations. Emerging work on “Visualization in Motion” [94] and “First-Person View Visualization” [34] provides valuable guidance in this direction. In addition, AR systems should support smooth transitions between embedded visualizations and traditional 2D visualizations in the headset depending on task context and user needs.
@@
Future Studies. In our study, participants selected targets verbally. A natural follow-up would involve embodied selection, e.g., walking to or pointing at targets, to examine whether the spatial support of AR-embedded visualizations becomes more critical under physical interaction. We hypothesize that in such embodied tasks, participants using 2D maps will incur greater cognitive load in spatial mapping, possibly diminishing their ability to evaluate targets effectively. Future work should explore how task modality (e.g., static vs. embodied) interacts with visualization and AI support in complex environments.
5.4 AR-Embedded Visualizations May Lower Cognitive Engagement Compared with 2D Minimap
@@
In our study, there was no significant difference in response time between Minimap+NoAI and AR+NoAI once initial search time was excluded. However, participants in AR+AI made decisions more quickly yet achieved lower overall accuracy than those in Minimap+AI. This pattern aligns with reliance on heuristic judgments – often described as Type 1 thinking [12, 22]. Together with the high self-reported confidence, these results suggest a lack of deeper cognitive engagement: once AI assistance was available, participants quickly felt confident enough in their intuitive judgments, rather than engaging in more analytical (Type 2) thinking. These findings imply that AR-embedded visualizations, by offering a highly immersive or embodied experience, may inadvertently reduce users’ motivation to scrutinize information.
@@
Design Implications. It remains unclear whether this phenomenon is specific to our task design, a broader characteristic of AR see-through interfaces, or a more general cognitive shift from 2D to AR-embedded. Design strategies that prompt more deliberation – such as “cognitive forcing functions” [10] – could be incorporated to encourage deeper analytic thinking in AR environments.
@@
Future Studies. Further research is needed to determine whether AR-embedded visualizations consistently reduces cognitive motivation compared to 2D visualizations, and to explore methods for mitigating AR-induced cognitive shortcuts across diverse tasks, visualization designs, and populations.
5.5 Study Limitations
@@
Our study offers an initial examination of AI-aided spatial decision-making with AR see-through visualizations. Although guided by prior research on AR visual systems for spatial tasks [91, 90], we acknowledge that our system design does not represent a fully optimized or universal solution. To ensure experimental control, we adopted fixed parameters, such as a discrete set of spatial arrangements, an AI accuracy of 75%, and a 20-second decision window, that may not capture the full complexity of high-stakes, real-world environments or more advanced AI systems. Consequently, our findings are most transferable to tasks resembling the indoor, time-sensitive context described in Section 2.1.
@@
Moreover, our participant pool primarily comprised well-educated individuals, which may not reflect the behavior of specialized user groups (e.g., emergency responders or security professionals). Finally, as is common in spatial cognition research [75], our sample size was modest; although sufficient for controlled experimentation, a larger and more diverse sample would strengthen external validity. We envision this work as a stepping stone for future studies to extend and refine AR-based decision support across broader tasks and populations.
6 Conclusion

### v2

@@ 5.2.1. Diverse Decision Strategies Across Visualizations
Summary. Compared to the Minimap conditions, participants in the X-ray conditions showed a stronger preference for spatial reasoning (distance) in their decision-making. This suggests that the AR X-ray visualization inherently facilitates spatial perception, yet they also function as a double-edged sword by introducing perceptual biases and occlusion challenges.
5.2.2. Increased Trust in AI Suggestions with X-Ray
The visualizations also impacted participants’ subjective trust on AI.
@@ 5.2.2. Increased Trust in AI Suggestions with X-Ray
Minimap: Make Own Choice First, Then Consult AI. In the Minimap conditions, most participants (22 out of 32) preferred first to form their own judgment and then cross-check against the AI suggestion, for instance, “I use it just to double-check my decision” (P22). Participants reported feeling more confident in their own judgment when using the Minimap visualization because it presented the information required to make informed choices. They typically consulted the AI in cases of uncertainty or indecision: “If I felt stuck between two locations, I would check the AI to break the tie” (P20).
@@
X-ray: AI as a Starting Anchor. Participants expressed greater trust in AI suggestions in the X-ray conditions, frequently using the AI suggestion as an initial reference or anchor for their decisions. This subjective feedback may help explain the observed over-reliance. Several participants such as P6 and P11 mentioned that AI helped narrow the decision space and enabled quicker comparisons. Participants provided several reasons for trusting AI more with X-ray:
@@
  • Perceptual and Navigation Challenges. Participants found distance estimation (e.g., P2), queue length assessment (e.g., P4, P5), and locating navigation paths (e.g., P12) more effortful in AR. Thus, AI suggestions offered a useful cognitive shortcut, aligning with findings from prior research on AI-aided decision-making (2021BansalDoesTheWholeExceedItsParts).
@@
  • Unfamiliarity with AR X-ray. Some participants like P31 mentioned that unfamiliarity with the AR X-ray increased their reliance on AI guidance.
@@
  • Enhanced Immersive Realism and Embodiment of AI suggestions. Interestingly, several participants such as P4 and P13, indicated that the realistic, immersive, embodied presentation of AI suggestions within the AR X-ray heightened their sense of trust and acceptance, which sometimes led to over-reliance on wrong AI suggestions.
@@
Summary. Compared to the Minimap, the X-ray visualization led to greater participant trust in and reliance on AI suggestions. Participants attributed this shift to greater perceptual complexity, unfamiliarity with the AR X-ray visualization, and the more realistic presentation of AI.
6. Discussion
@@
6.1. Perceptual Challenges of AR X-Ray Can Lead to Over-Reliance on AI
Our findings suggest that perceptual challenges of the AR X-ray, such as occlusion, large canvas space, and spatial estimation difficulty, can increase users’ reliance on AI suggestions, even when those suggestions are suboptimal. This over-reliance does not appear to stem from blind trust in AI, but rather from the inherent complexity of perceiving and evaluating visual information in large-scale physical environments. While prior research has discussed AR’s perceptual limitations (Drascic1996PerceptualII), our study directly connects these challenges to AI reliance patterns in time-sensitive spatial decision-making. This highlights the need for better embedded visualization designs that support spatial reasoning, especially in large-scale environments.
Design Implications. Visually embedding AI support in physical environments is not inherently beneficial. It can be neutral or even detrimental when perceptual challenges make it harder to verify AI suggestions. These findings underscore the need to rethink how AI assistance is integrated into AR, rather than directly porting AI designs from desktop or 2D interfaces. Designers should explicitly consider how spatial perception difficulties affect users’ ability to evaluate AI output.
Future Studies. In our study, situational data like queue length was visualized using human-shaped glyphs. A natural possibility is to compare visual encoding, for example, glyphs vs. numeric labels, to evaluate how data representation (e.g., legibility, uncertainty, and saliency) influences AI reliance.
Additionally, examining how different fundamental characterizations of the situation (e.g., task difficulty, information ambiguity, time pressure and many others) modulate trust and reliance in AR embedded visualizations would offer deeper insight into AR-AI interaction design.
6.2. Cognitive Biases Triggered by AR X-Ray Can Induce Inappropriate AI Reliance
Beyond perceptual difficulty, our qualitative findings point to cognitive biases that are unique to the AR X-ray. Participants reported visual proximity illusions and increased trust in AI suggestions due to their spatial embodiment in the environment. These effects align with prior research on spatial cognition biases (raghubir1996crow) and trust in embodied agents (Kim2018DigitalAssistantNeedABody).
Although cognitive biases have been studied in traditional visualization contexts (dimara2018task; wall2021left), their implications in embodied, spatial, and AI-augmented AR systems remain underexplored. As Padilla et al. (padilla2018decision) suggested, when systems are designed with such biases in mind, users can benefit from Type 1 (intuitive) thinking—particularly under time pressure. Our findings highlight a fertile area for interdisciplinary research at the intersection of spatial cognition, embedded visualizations, and human-AI decision-making.
@@ 6.2. Cognitive Biases Triggered by AR X-Ray Can Induce Inappropriate AI Reliance
6.3. Aligning the Strengths of AR Embedded Visualizations with Spatial Tasks
@@
In our study, while the X-ray did not outperform the Minimap on decision accuracy, it significantly reduced post-trial pointing errors, suggesting improved spatial mapping. This distinction suggest that AR embedded visualizations may be more beneficial in continuous, action-based spatial workflows rather than in isolated decision-making tasks.
@@
In spatial tasks like emergency response, selecting a correct location is only the first step; what follows is action. Misinterpreting a target’s location after selection can be consequential. For instance, a user might correctly choose an exit from a wall-mounted map but walk the wrong way due to reference-frame confusion. AR embedded visualizations mitigate this risk by spatially anchoring information in the environment. Thus, rather than emphasizing the weakness of AR embedded visualizations, our findings illustrate their potential strengths in embodied spatial tasks and highlight when and where AR embedded visualizations are best applied.
@@
Design Implications. AR embedded visualizations may be most effective when tightly coupled with tasks involving physical action, rather than static decision-making. Designers should consider the viewer’s movement and real-time orientation when designing the embedded visualizations. Emerging work on “Visualization in Motion” (DBLP:journals/IEEE_TVCG/YaoBVI22) and “First-Person View Visualization” (isenberg:hal-04782928) provides valuable guidance in this direction. In addition, AR systems should support smooth transitions between embedded visualizations and traditional 2D visualizations in the headset depending on task context and user needs.
@@
Future Studies. In our study, participants selected targets verbally. A natural follow-up would involve embodied selection, for example, walking to or pointing at targets, to examine whether the spatial support of AR embedded visualizations becomes more critical under physical interaction. We hypothesize that in such embodied tasks, participants using 2D maps will incur greater cognitive load in spatial mapping, possibly diminishing their ability to evaluate targets effectively. Future work should explore how task modality (e.g., static vs. embodied) interacts with visualization and AI support in complex environments.
6.4. AR X-Ray May Lower Cognitive Engagement Compared with 2D Visualizations
@@
In our study, there was no significant difference in response time between the Minimap+NoAI and X-ray+NoAI conditions once initial search time was excluded. However, participants in the X-ray+AI condition made decisions more quickly yet achieved lower overall accuracy than those in the Minimap+AI condition. This pattern aligns with reliance on heuristic judgments, often described as Type 1 thinking (Chaiklin2012ThinkingFA; JSBTDualProcessTheory2013). Together with the high self-reported confidence, these results suggest a lack of deeper cognitive engagement: once AI assistance was available, participants quickly felt confident enough in their intuitive judgments, rather than engaging in more analytical (Type 2) thinking. These findings imply that AR embedded visualizations, by offering a highly immersive or embodied experience, may inadvertently reduce users’ motivation to scrutinize information.
@@
Design Implications. It remains unclear whether this lower cognitive engagement is specific to our task design, a broader characteristic of AR embedded visualizations, or a more general cognitive shift from 2D to AR egocentric perspective. Design strategies that prompt more deliberation, such as “cognitive forcing functions” (2021ToTrustOrToThink), could be incorporated to encourage deeper analytic thinking in AR environments.
@@
Future Studies. Further research is needed to determine whether AR embedded visualizations consistently reduce cognitive motivation compared to 2D visualizations, and to explore methods for mitigating AR-induced cognitive shortcuts across diverse tasks, visualization designs, and populations.
6.5. Study Limitations
@@
Our study offers an initial examination of AI-aided spatial decision-making with AR embedded visualizations. Although guided by prior research on AR visual systems for spatial tasks (XUImroveIndoorWayfinding; XUARteambasedsearch), we acknowledge that our system design does not represent a fully optimized, advanced, or universal solution.
@@
To ensure experimental control, we adopted fixed parameters, such as a discrete set of spatial arrangements, an AI accuracy of 75% (see Section 4.7), and a 20-second decision window, which may not capture the full complexity of high-stakes, real-world environments or more advanced AI systems. Consequently, our findings are most transferable to tasks resembling the indoor, time-sensitive context described in Section 3.1. Moreover, our participant pool primarily comprised well-educated individuals, which may not reflect the behavior of specialized user groups, for example, emergency responders or security professionals. As is common in spatial cognition research (Tan2024InvisibleME), our sample size was modest; although sufficient for controlled experimentation, a larger and more diverse sample would strengthen external validity.
Finally, in our analysis, there may be a minor initial search time in the Minimap condition, which depends on how quickly participants detect the first two targets in the Minimap. However, this subtle interval cannot be reliably measured from first-person video alone. Future work incorporating eye-tracking data, which is currently unavailable on Apple Vision Pro, could help capture this duration more accurately. We also acknowledge that during S2, some cognitive evaluation of the first target may occur. Nevertheless, by definition, a decision can only be made once the second target has been located. Overall, we view this study as a scientifically grounded first step that provides a reasonable and well-documented foundation for subsequent research to extend and refine AR-based decision support across broader tasks and populations.
7. Conclusion